### v2

# Some Plancherel identities for unbounded subsets of $\mathbb{R}$ in duality

Abstract.
In relation to Fuglede’s conjecture, we establish several Plancherel-type identities and demonstrate the surjectivity of the Fourier transform between certain unbounded tiling sets of $\mathbb{R}$ that are in duality. In the terminology commonly used in the context of Fuglede’s conjecture, our result states that an open set tiles $\mathbb{R}$ by the finite set $\{0,1,\dots,p-1\}$ if and only if it admits a spectrum (or, equivalently, a dual pair measure) given by the Lebesgue measure on $\left[-\tfrac{1}{2p},\tfrac{1}{2p}\right]+\mathbb{Z}$.
Key words and phrases:
differential operator, self-adjoint operator, Fourier bases, Fourier transform, Plancherel identity, Fuglede conjecture, lattice, tile
2010 Mathematics Subject Classification:
47E05,42A16
1. Introduction
In 1958, Irving Segal posed the following question to Bent Fuglede: Let $\Omega$ be an open subset of $\mathbb{R}^{d}$. Consider the partial differential operators
$D_{1}=\frac{1}{2\pi i}\frac{\partial}{\partial x_{1}},\dots,D_{d}=\frac{1}{2\pi i}\frac{\partial}{\partial x_{d}},$
defined on the space $C_{0}^{\infty}(\Omega)$ of smooth, compactly supported functions on $\Omega$. Under what conditions do these differential operators admit commuting (unbounded) self-adjoint extensions $H_{1},\dots,H_{d}$ on $L^{2}(\Omega)$, where commutation is understood in the sense of their spectral measures?
In his seminal 1974 paper [FUG74], Fuglede provided an answer to Segal’s question in the case where $\Omega$ is a connected, finite-measure, Nikodym domain—that is, when the Poincaré inequality holds:
Theorem 1.1.
[FUG74, Theorem I] Let $\Omega\subset\mathbb{R}^{d}$ be an open, connected Nikodym region of finite measure. Define
$e_{\lambda}(x)=e^{2\pi i\lambda\cdot x},\quad(x\in\mathbb{R}^{d},\lambda\in\mathbb{R}^{d}).$
There exist commuting self-adjoint extensions of the differential operators $\{D_{j}:j=1,\dots,d\}$ if and only if there exists a subset $\Lambda$ of $\mathbb{R}^{d}$ such that the family of exponential functions
$\{e_{\lambda}:\lambda\in\Lambda\}$
forms an orthogonal basis for $L^{2}(\Omega)$.
Definition 1.2.
A measurable set $\Omega\subset\mathbb{R}^{d}$ of finite Lebesgue measure is called spectral if there exists a set $\Lambda\subset\mathbb{R}^{d}$ such that the family of exponential functions
$\{e_{\lambda}:\lambda\in\Lambda\}$
forms an orthogonal basis for $L^{2}(\Omega)$. In this case, $\Lambda$ is called a spectrum for $\Omega$.
We say that $\Omega$ tiles $\mathbb{R}^{d}$ by translations if there exists a set $\mathcal{T}\subset\mathbb{R}^{d}$ such that the translates $\{\Omega+t:t\in\mathcal{T}\}$ form a partition of $\mathbb{R}^{d}$ up to measure zero. The set $\mathcal{T}$ is then called a tiling set for $\Omega$, and we say that $\Omega$ tiles $\mathbb{R}^{d}$ by $\mathcal{T}$.
Thus, in this framework, when $\Omega$ is connected, has finite measure, and is a Nikodym domain, there exist commuting self-adjoint extensions of the differential operators $\{D_{j}\}$ if and only if $\Omega$ is spectral. Because this characterization is somewhat abstract, Fuglede proposed his famous conjecture:
Conjecture 1.3.
[Fuglede’s Conjecture] A measurable subset $\Omega\subset\mathbb{R}^{d}$ of finite Lebesgue measure is spectral if and only if it tiles $\mathbb{R}^{d}$ by translations.
The conjecture was later shown to be false in dimensions $d\geq 3$ [TAO04, FMM06], but it remains true for convex domains [LM22].
In [PED87], Steen Pedersen extended Fuglede’s result by removing both the Nikodym condition and the finite-measure assumption. Naturally, the definition of a spectral set had to be adapted for infinite-measure domains (since $e_{\lambda}$ is not square-integrable in that case).
Definition 1.4.
For a function $f\in L^{1}(\mathbb{R}^{d})$, define its classical Fourier transform as
$\hat{f}(x)=\int_{\mathbb{R}^{d}}f(x)e^{-2\pi it\cdot x}\,dx,\quad(t\in\mathbb{R}^{d}).$
Let $\Omega\subset\mathbb{R}^{d}$ be measurable and let $\mu$ be a positive Radon measure on $\mathbb{R}^{d}$. We say that $(\Omega,\mu)$ is a spectral pair if: (1) for each $f\in L^{1}(\Omega)\cap L^{2}(\Omega)$, the continuous function $t\mapsto\hat{f}(t)$ satisfies $\int|\hat{f}|^{2}\,d\mu<\infty$; and (2) the mapping $f\mapsto\hat{f}$ from $L^{1}(\Omega)\cap L^{2}(\Omega)\subset L^{2}(\Omega)$ into $L^{2}(\mu)$ is isometric and has dense range.
This map then extends by continuity to an isometric isomorphism
$\mathscr{F}:L^{2}(\Omega)\to L^{2}(\mu).$
The isometry property means that we have a Plancherel identity implemented by taking the Fourier transform of functions in $L^{2}(\Omega)$ and restricting it to the support of the measure $\mu$ and integrating the absolute value squared of this restriction of the Fourier transform against the measure $\mu$.
The set $\Omega$ is called spectral if there exists a measure $\mu$ such that $(\Omega,\mu)$ forms a spectral pair; $\mu$ is then called a pair measure or a dual measure for $\Omega$. When $\Omega$ has finite measure, this definition coincides with the earlier one [PED87, Corollary 1.11], and the dual measure $\mu$ is the counting measure on the spectrum $\Lambda$.
With this broader definition, Pedersen generalized Fuglede’s theorem by eliminating both the Nikodym and finite-measure conditions, while maintaining the assumption that $\Omega$ is connected.
Theorem 1.5.
[PED87, Theorem 2.2] Let $\Omega$ be an open, connected subset of $\mathbb{R}^{d}$. Then the partial differential operators $\{D_{j}\}$ admit commuting self-adjoint extensions if and only if $\Omega$ is a spectral set.
The results were then extended to disconnected domains in [CD25]. In this case, if $\Omega$ is spectral, then there are commuting self-adjoint extensions of the differential operators $\{D_{j}\}$. However, the converse has to be adjusted a little: the Fourier transform in Definition 1.4 is allowed to have some weights which differ on the components of $\Omega$, and the resulting condition is necessary and sufficient.
In the 1974 paper Fuglede proved that his conjecture is true when the spectrum or the tiling set is a lattice in $\mathbb{R}^{d}$.
Theorem 1.6.
Let $\Omega$ be a measurable subset of $\mathbb{R}^{d}$ of finite measure and let $A$ be some invertible $d\times d$ matrix. Then $\Omega$ tiles $\mathbb{R}^{d}$ by the lattice $A\mathbb{Z}^{d}$ if and only if $\Omega$ has spectrum the dual lattice $(A^{T})^{-1}\mathbb{Z}^{d}$.
Many examples of spectral sets are known in the case when $\Omega$ has finite measure, and the spectrum can be a non-lattice, see e.g, [KOL24]. However, much less is known for sets of infinite measure. Of course, the simplest, and most important example of a spectral set of infinite measure is $\mathbb{R}^{d}$, with pair measure the Lebesgue measure on $\mathbb{R}^{d}$. To obtain some other examples, a result similar to Theorem 1.6 was proved for sublattices of $\mathbb{R}^{d}$ in [CD25]:
Theorem 1.7.
Write $\mathbb{R}^{d}=\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}$. Let $A$ be an invertible real $d\times d$ matrix. The set $\Omega$ tiles with the discrete subgroup $\mathcal{T}=A(\mathbb{Z}^{d_{1}}\times\{0\})$ if and only if $\Omega$ has pair measure $A^{T}({}_{\mathbb{Z}^{d_{1}}}\times\mathfrak{m}_{d_{2}})$ which is supported on the dual set $\mathcal{T}^{*}$,
[…]
Here the Dirac comb $_{\mathbb{Z}^{d_{1}}}$ is the counting measure on $\mathbb{Z}^{d_{1}}$, and $\mathfrak{m}_{d_{2}}$ is the Lebesgue measure on $\mathbb{R}^{d_{2}}$.
For a Borel measure $\mu$ on $\mathbb{R}^{d}$, the measure $A^{T}\mu$ is defined by $A^{T}\mu(E)=|\det(A^{T})^{-1}|\mu(A^{T}E)$ for any Borel subset $E$ in $\mathbb{R}^{d}$. Equivalently, for any compactly supported continuous function $f$ on $\mathbb{R}^{d}$:
(1.1) $\int f\,dA^{T}\mu=|\det(A^{T})^{-1}|\int f((A^{T})^{-1}x)\,d\mu(x).$
Note that if the measure $\mu$ is supported on a set $\Lambda$, then $A^{T}\mu$ is supported on $(A^{T})^{-1}\Lambda$.
In this paper we provide a new class of examples of unbounded spectral sets in $\mathbb{R}$ and we prove the following theorem:
Theorem 1.8.
Let $\Omega$ be an open subset in $\mathbb{R}$ and let $p\in\mathbb{N}$, $p\geq 2$. Then $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$ if and only if $\Omega$ is a spectral set with pair measure $\mu$ equal to a renormalized Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, $\mu=p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$ .
2. Proof of Theorem 1.8
We will need some definitions and notations.
Definition 2.1.
Let $\Omega_{0}$ and $\Omega$ be two measurable subsets in $\mathbb{R}$, and let $\mathcal{T}$ be a finite or countable subset of $\mathbb{R}$. We say that $\Omega_{0}$ tiles $\Omega$ by $\mathcal{T}$ if $\{\Omega_{0}+t:t\in\mathcal{T}\}$ is a partition of $\Omega$, up to measure zero.
We denote by $T_{a}$ the translation operator on functions on $\mathbb{R}$
$(T_{a})f(x)=f(x-a),\quad(x\in\mathbb{R},a\in\mathbb{R}).$
Definition 2.2.
Let $E$ and $F$ be two measurable subsets of $\mathbb{R}$. We say that $E$ and $F$ are congruent $\operatorname{mod}\mathbb{Z}$ if there exists a partition (up to measure zero) of $E$ into possibly empty sets $\{E_{k}:k\in\mathbb{Z}\}$ such that $\{E_{k}+k:k\in\mathbb{Z}\}$ is a partition of $F$ (up to measure zero).
Definition 2.3.
Let $\mu$ be a finite Radon measure on $\mathbb{R}$ and $\Lambda\subset\mathbb{R}$. We say that $\mu$ is a spectral measure with spectrum $\Lambda$ if the family $\{e_{\lambda}:\lambda\in\Lambda\}$ forms an orthogonal basis for $L^{2}(\mu)$.
Definition 2.4.
For a measurable subset $E$ of $\mathbb{R}$, we denote by $|E|$ its Lebesgue measure and by $\chi_{E}$ its characteristic function. For a finite subset $F$, we denote by $|F|$ its cardinality. We denote by $\delta_{\gamma}$ the Dirac measure at the point $\gamma$ in $\mathbb{R}$. For a discrete subset $\Gamma$ of $\mathbb{R}$, we define the Dirac comb on $\Gamma$ to be the counting measure on $\Gamma$,
$_{\Gamma}=\sum_{\gamma\in\Gamma}\delta_{\gamma}.$
For a measurable subset $E$ of $\mathbb{R}$, we denote by $\mbox{Leb}_{E}$ the Lebesgue measure restricted to $E$.
For a function $f$ on $\mathbb{R}$ we denote its periodization by
$\operatorname*{Per}(f)(t)=\sum_{k\in\mathbb{Z}}f(t+k),\quad(t\in\mathbb{R}).$
We begin now the proof of our main result.
Proof of Theorem 1.8.
We assume that $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$.
Proposition 2.5.
Suppose $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$. Define $\Omega_{0}=\Omega\cap[0,p)$.
  1. (i)
    $\Omega+kp\subset\Omega$ for all $k\in\mathbb{Z}$.
  2. (ii)
    $\Omega_{0}$ tiles $\Omega$ by $p\mathbb{Z}$; in particular
    $\Omega=\Omega_{0}+p\mathbb{Z}.$
  3. (iii)
    $\Omega_{0}$ tiles $\mathbb{R}$ by $\mathbb{Z}$. Therefore $\Omega_{0}$ has measure 1, has spectrum $\mathbb{Z}$, and is congruent to $[0,1)$ modulo $\mathbb{Z}$.
Proof.
Let $x\in\Omega$. Then $x+p\in\Omega$. Indeed, since $\Omega$ tiles with $\{0,\dots,p-1\}$ we can write $x+p=y+j$ for $y\in\Omega$ and $j\in\{0,\dots,p-1\}$. If $j\neq 0$, then x+(pj)=yx+(p-j)=y so $y\in\Omega\cap(\Omega+(p-j))$, which can happen only for $y$ in a measure zero set.
If xp=y+jx-p=y+j with $x\in\Omega$ and $j\in\{0,1,\dots,p-1\}$ then $\Omega\ni x=y+p+j\in\Omega+j$ and therefore $j$ has to be 0 and $x-p\in\Omega$. By induction, we get $x+pk\in\Omega$, for $k\in\mathbb{Z}$. This proves (i).
For (ii), it is clear from (i) that $\Omega_{0}+p\mathbb{Z}\subseteq\Omega$. Now take $x\in\Omega$. Then we can write $x=y+pk$ for some $y\in[0,p)$ and $k\in\mathbb{Z}$. With (i), $y=x-kp\in\Omega$. Thus $y\in\Omega_{0}$ and $x\in\Omega_{0}+p\mathbb{Z}$. This proves the reverse inclusion.
Since $\Omega_{0}$ is contained in $[0,p)$, the sets $\Omega_{0}+pk$, $k\in\mathbb{Z}$, are mutually disjoint. This proves (ii).
For (iii), note that
$\Omega_{0}+\mathbb{Z}=\Omega_{0}+p\mathbb{Z}+\{0,\dots,p-1\}=\Omega+\{0,\dots,p-1\}=\mathbb{R}.$
The sets $\Omega_{0}+pk+j$, with $k\in\mathbb{Z}$ and $j\in\{0,\dots,p-1\}$ are disjoint by (ii) and the fact that $\Omega$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$. Thus $\Omega_{0}$ tiles $\mathbb{R}$ by $\mathbb{Z}$.
With Theorem 1.6, we get that $\Omega_{0}$ has spectrum $\mathbb{Z}$. Also, let $\Omega_{0,k}=\Omega_{0}\cap[k,k+1)$ and let $I_{k}=\Omega_{0,k}-k\subset[0,1)$, for all $k\in\mathbb{Z}$. Then the sets $\{I_{k}:k\in\mathbb{Z}\}$ are disjoint (otherwise the sets $\Omega_{0}+k$ are not), and their union is $[0,1)$ because, for $x\in[0,1)$ there is a unique $k\in\mathbb{Z}$ and $x_{0}\in\Omega_{0}$ such that $x=x_{0}-k$. So $x_{0}\in\Omega_{0,k}$ and $x\in I_{k}$. Thus $\Omega_{0}$ is congruent to $[0,1)$ modulo $\mathbb{Z}$ and
$|\Omega_{0}|=\sum_{k\in\mathbb{Z}}|\Omega_{0,k}|=\sum_{k\in\mathbb{Z}}|I_{k}|=1.$
Proposition 2.6.
For $n\in\mathbb{N}$ let
$\Omega_{n}:=\Omega_{0}+p\{-n,-n+1,\dots,n-1,n\}.$
  1. (i)
    The sets $\Omega_{n}$ are increasing and cover $\Omega$:
    $\cup_{n\in\mathbb{N}}\Omega_{n}=\Omega.$
  2. (ii)
    The set $\Omega_{n}$ tiles $\Omega$ by $p(2n+1)\mathbb{Z}$.
  3. (iii)
    The set $\Omega_{n}$ has spectrum
    $\Lambda_{n}:=\frac{1}{p(2n+1)}\left\{-n,-n+1,\dots,n\right\}+\mathbb{Z}.$
Proof.
(i) is clear, from Proposition 2.5(ii). For (ii), note that the set $L_{n}=\{-n,\dots,n\}$ tiles $\mathbb{Z}$ by $(2n+1)\mathbb{Z}$. Therefore
$\Omega_{n}+p(2n+1)\mathbb{Z}=\Omega_{0}+pL_{n}+p(2n+1)\mathbb{Z}=\Omega_{0}+p\mathbb{Z},$
and the sets $\Omega_{n}+p(2n+1)k=\Omega_{0}+pL_{n}+p(2n+1)k$, $k\in\mathbb{Z}$ are disjoint.
For (iii), we will use the following well known lemma (see e.g., [JP98]). We include the proof for convenience.
Lemma 2.7.
Let $\mu$ be finite Radon measure on $\mathbb{R}$ with total measure $\mu(\mathbb{R})=m$, and let $\Lambda$ be a finite or countable subset of $\mathbb{R}$. The measure $\mu$ has spectrum $\Lambda$ if and only if
(2.1) $\sum_{\lambda\in\Lambda}|\hat{\mu}(t+\lambda)|^{2}=m^{2},$
where $\hat{\mu}$ is the Fourier transform of the measure $\mu$,
$\hat{\mu}(t)=\int e^{-2\pi itx}\,d\mu(x),\quad(t\in\mathbb{R}).$
Proof.
Suppose $\Omega$ has spectrum $\Lambda$. Then
$\{\frac{1}{\sqrt{m}}e_{\lambda}:\lambda\in\Lambda\}$
is an orthonormal basis for $L^{2}(\mu)$. Applying the Parseval identity to the function $e_{-t}$, $t\in\mathbb{R}$ we obtain
$m=\|e_{-t}\|^{2}=\sum_{\lambda\in\Lambda}\left|\left\langle e_{-t}\,,\,\frac{1}{\sqrt{m}}e_{\lambda}\right\rangle_{L^{2}(\mu)}\right|^{2}=\frac{1}{m}\sum_{\lambda\in\Lambda}\left|\hat{\mu}(t+\lambda)\right|^{2}$
and this implies (2.1).
For the converse, if we plug $t=-\lambda_{0}$ in (2.1), since $\hat{\mu}(0)=|\Omega|=m$, we get that $\hat{\mu}(-\lambda_{0}+\lambda)=0$ for $\lambda\neq\lambda_{0}$, and this proves that the exponentials $e_{\lambda}$ and $e_{\lambda_{0}}$ are orthogonal. With the same computation as above, (2.1) then shows that
$\|e_{-t}\|^{2}=\sum_{\lambda\in\Lambda}\left|\left\langle e_{-t}\,,\,\frac{1}{\sqrt{m}}e_{\lambda}\right\rangle_{L^{2}(\mu)}\right|^{2}.$
On the right hand side we have the norm squared of the projection of $e_{-t}$ onto the closed span of the functions $\{e_{\lambda}:\lambda\in\Lambda\}$. This shows that $e_{-t}$ is in this closed span. By the Stone-Weierstrass theorem and by approximation by continuous compactly supported functions, the functions $e_{-t}$ span the entire space $L^{2}(\mu)$. It follows that $\{e_{\lambda}:\lambda\in\Lambda\}$ also span the entire space $L^{2}(\mu)$ and therefore they form an orthogonal basis for it. ∎
Returning to the proof of Proposition 2.6(iii), we apply Lemma 2.7 to the Lebesgue measure on $\Omega_{n}$. Note that this measure is the convolution $\mbox{Leb}_{\Omega_{n}}=\mbox{Leb}_{\Omega_{0}}\ast{}_{pL_{n}}$. Therefore its Fourier transform is
$\hat{\chi}_{\Omega_{n}}(t)=\hat{\chi}_{\Omega_{0}}(t)\hat{\Sha}_{pL_{n}}(t).$
We also have
$\hat{\Sha}_{pL_{n}}(t)=\int e^{-2\pi itx}\,d{}_{pL_{n}}(x)=\sum_{l=-n}^{n}e^{-2\pi iplt},$
which implies that $\hat{\Sha}_{pL_{n}}$ has period $1/p$.
It is also easy to check that the measure $_{pL_{n}}$ has spectrum $\frac{1}{p(2n+1)}L_{n}$ (this is basically the discrete Fourier transform on $\mathbb{Z}_{2n+1}$, rescaled by $p$). Therefore, with Lemma 2.7,
(2.2) $\sum_{j\in L_{n}}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}\right)\right|^{2}=|L_{n}|^{2}=(2n+1)^{2},\quad(t\in\mathbb{R}).$
Since $\Omega_{0}$ has spectrum $\mathbb{Z}$, we have that
(2.3) $\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{0}}(t+k)\right|^{2}=|\Omega_{0}|^{2}=1,\quad(t\in\mathbb{R}).$
We compute
$\sum_{j\in L_{n}}\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{n}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}$
$\sum_{j\in L_{n}}\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{0}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}$
and using the periodicity of $\hat{\Sha}_{pL_{n}}$,
$\sum_{j\in L_{n}}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}\right)\right|^{2}\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{0}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}$
and with (2.3), and then (2.2),
$\sum_{j\in L_{n}}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}\right)\right|^{2}\cdot 1=(2n+1)^{2}=|\Omega_{n}|^{2}.$
Then, with Lemma 2.7, we conclude that $\frac{1}{p(2n+1)}L_{n}+\mathbb{Z}$ is a spectrum for $\Omega_{n}$. ∎
Next we prove the first condition for $\Omega$ to be spectral, from Definition 1.4, namely that the Fourier transform is isometric from $L^{2}(\Omega)$ to the $L^{2}$-space of the pair measure, which is $p$ times the Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$. The notation for this $L^{2}$-space is $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx\right)$.
Proposition 2.8.
The Fourier transform
$L^{2}(\Omega)\ni f\to\hat{f}|_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}\in L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx\right)$
is an isometry.
Proof.
Let $f$ be a $C^{\infty}$-function compactly supported in $\Omega$. Then $\hat{f}$ is a Schwartz function.
Let $n$ be big enough so that the support of $f$ is covered by $\Omega_{n}=\Omega_{0}+p\{-n,\dots,n\}$. Since $\Omega_{n}$ has spectrum $\frac{1}{p(2n+1)}L_{n}+\mathbb{Z}$, we have, using the Parseval identity,
$\int_{\Omega}|f(x)|^{2}\,dx=\int_{\Omega_{n}}|f(x)|^{2}\,dx=\sum_{j\in L_{n},k\in\mathbb{Z}}\left|\left\langle f\,,\,\frac{1}{\sqrt{2n+1}}e_{\frac{j}{p(2n+1)}+k}\right\rangle_{L^{2}(\Omega_{n})}\right|^{2}$
$\frac{1}{2n+1}\sum_{j\in L_{n},k\in\mathbb{Z}}\left|\hat{f}\left(\frac{j}{p(2n+1)}+k\right)\right|^{2}.$
This is a Riemann sum for the function $|\hat{f}|^{2}$ with sampling points $\frac{j}{p(2n+1)}+k$, $j\in L_{n}$, $k\in\mathbb{Z}$, and thus the length of the intervals of the partition is $\frac{1}{p(2n+1)}$. Since the common factor is $\frac{1}{2n+1}$ we have to adjust it by the constant $p$, and therefore the function that is integrated is actually $p|\hat{f}|^{2}$ on the set $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$.
Since $\hat{f}$ is a Schwartz function, these Riemann sums converge to the integral
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}|\hat{f}(t)|^{2}p\,dt.$
This implies that
$\int_{\Omega}|f(x)|^{2}\,dx=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}|\hat{f}(t)|^{2}p\,dt.$
Since $C_{0}^{\infty}(\Omega)$ is dense in $L^{2}(\Omega)$, we get that the Fourier transform establishes an isometry between the two $L^{2}$-spaces.
Proposition 2.9.
The Fourier transform from $L^{2}(\Omega)$ to $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx\right)$ is onto.
Proof.
First, note that since $\Omega_{0}$ is congruent to $[0,1]$ $\operatorname{mod}\mathbb{Z}$, we have for $k\in\mathbb{Z}$,
(2.4) $\hat{\chi}_{\Omega_{0}}(k)=\int_{\Omega_{0}}e^{-2\pi ikx}\,dx=\int_{[0,1]}e^{-2\pi ikx}\,dx=\left\{\begin{array}[]{cc}1,&k=0,\\ 0,&k\neq 0.\end{array}\right$.
The function $\chi_{\Omega}=\chi_{\Omega_{0}+p\mathbb{Z}}$ has period $p$. Therefore, its Fourier transform, as a tempered distribution, is
(2.5) $\hat{\chi}_{\Omega}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}.$
This follows from the next lemma:
Lemma 2.10.
Let $\tilde{h}$ be a function of period $p$ on $\mathbb{R}$ and $h=\tilde{h}|_{[0,p]}\in L^{2}[0,p]$. Represent $\tilde{h}$ in the Fourier basis $\frac{1}{\sqrt{p}}e_{k/p}$, $k\in\mathbb{Z}$:
$\tilde{h}=\sum_{k\in\mathbb{Z}}\frac{1}{p}\left\langle h\,,\,e_{k/p}\right\rangle_{L^{2}[0,p]}e_{k/p}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)e_{k/p}.$
Then the Fourier transform of $\tilde{h}$ as a tempered distribution is
(2.6) $\hat{\tilde{h}}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)\delta_{k/p}.$
Proof.
Let $\Lambda_{\tilde{h}}$ be the tempered distribution associated to $\tilde{h}$,
$\Lambda_{\tilde{h}}(\varphi)=\int_{\mathbb{R}}\tilde{h}(x)\varphi(x)\,dx,\quad(\varphi\mbox{ Schwartz function}).$
Then, for a Schwartz function $\varphi$,
$\hat{\Lambda}_{\tilde{h}}(\varphi)=\Lambda_{\tilde{h}}(\hat{\varphi})=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)\int_{\mathbb{R}}e^{2\pi itk/p}\hat{\varphi}(t)\,dt$
with the Fourier Inversion Formula,
$\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)\varphi(k/p),$
and (2.6) follows. The sums are convergent because $\sum|\hat{h}(k/p)|^{2}<\infty$ and $\varphi,\hat{\varphi}$ are Schwartz functions, thus they have rapid decay.
Take now some $C_{0}^{\infty}$-function $g$ supported on one of the intervals $\left[-\frac{1}{2p},\frac{1}{2p}\right]+k_{0}$ for some $k_{0}\in\mathbb{Z}$, and let $\varphi$ be its inverse Fourier transform; it is a Schwartz function. Let
$f=p\chi_{\Omega}\cdot\varphi.$
Then
$\hat{f}(t)=p\hat{\chi}_{\Omega}\ast\hat{\varphi}(t)=\left(\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}\right)\ast g(t)=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)g(t-k/p).$
We need to restrict $\hat{f}$ to $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$. Note that g(tk/p)g(t-k/p) is supported on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\frac{k}{p}+k_{0}$, and these sets are disjoint for different $k$’s. Thus, when we restrict $\hat{f}$ to $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, we will keep only the terms with $k=pl$ for $l\in\mathbb{Z}$:
$\hat{f}\cdot\chi_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}(t)=\sum_{l\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(pl/p)g(t-pl/p).$
But, by (2.4), $\hat{\chi}_{\Omega_{0}}(l)=0$ unless $l=0$, and therefore the only term left is $g(t)$. Thus
$\hat{f}|_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}=g.$
Since the linear combinations of functions like $g$ are dense in $L^{2}(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx)$ and since we proved that the map is an isometry, it follows that the map is onto. ∎
Propositions 2.8 and 2.9 show the direct implication: if $\Omega$ tiles with $\{0,1,\dots,p-1\}$ then it is spectral with the desired pair measure.
We move on now to the converse and assume that $\Omega$ is spectral with the given pair measure.
Proposition 2.11.
[…]
Proof.
Let $a\in\mathbb{R}$. We have
$\left\langle f\,,\,g\right\rangle_{L^{2}(\Omega)}=\left\langle e_{a}f\,,\,e_{a}g\right\rangle_{L^{2}(\Omega)}=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}\hat{f}(t-a)\overline{\hat{g}}(t-a)p\,dt$
$\sum_{k\in\mathbb{Z}}\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\hat{f}(t-a+k)\overline{\hat{g}}(t-a+k)p\,dt$
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t-a)p\,dt=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t)p\,dt.$
Let $h(t)=\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t)$. Then the function
$a\to\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a}h(t)\,dt$
is constant.
Let $a_{1}=\frac{1}{2p}+b+\epsilon$, $a_{2}=\frac{1}{2p}+b$, for $b\in\mathbb{R}$ and $\epsilon>0$. We have
$0=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a_{1}}h(t)\,dt-\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a_{2}}h(t)\,dt=\int_{[b+\epsilon,b+\frac{1}{p}+\epsilon]}h(t)\,dt-\int_{[b,b+\frac{1}{p}]}h(t)\,dt$
$\int_{[b+\frac{1}{p},b+\frac{1}{p}+\epsilon]}h(t)\,dt-\int_{[b,b+\epsilon]}h(t)\,dt.$
Divide by $\epsilon$ and, with the Lebesgue Differentiation Theorem, we get that for almost every $b\in\mathbb{R}$,
$h(b+1/p)=h(b),$
[…]
Proposition 2.12.
Suppose $\Omega$ is spectral with pair measure $p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$. Then $\Omega\cap(\Omega+j)=\emptyset$ up to measure zero, for all $j\in\mathbb{Z}$, $j\not\equiv 0\operatorname{mod}p$.
Proof.
Let $f,g\in L^{2}(\Omega)$. We compute, with Plancherel’s identity,
$\left\langle f\,,\,T_{j}g\right\rangle_{L^{2}(\mathbb{R})}=\int_{\mathbb{R}}{\hat{f}(t)}\overline{e^{-2\pi ijt}\hat{g}(t)}\,dt=\sum_{l\in\mathbb{Z}}\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\hat{f}(t+l/p)e^{2\pi ij(t+l/p)}\overline{\hat{g}}(t+l/p)\,dt$
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\sum_{i=0}^{p-1}\sum_{m\in\mathbb{Z}}\hat{f}(t+i/p+m)\overline{\hat{g}}(t+i/p+m)e^{2\pi ij(t+i/p)}\,dt$
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\sum_{i=0}^{p-1}e^{2\pi ij(t+i/p)}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t+i/p)\,dt$
with Propositon 2.11,
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t)e^{2\pi ijt}\sum_{i=0}^{p-1}e^{2\pi iji/p}\,dt=0,$
since $j\not\equiv 0(\operatorname{mod}p)$.
Suppose now that $\Omega\cap(\Omega+j)$ has positive measure. Then take $E\subset\Omega\cap(\Omega+j)$ of finite measure. Let $f=\chi_{E}$ and $g=\chi_{E-j}$. Then $f,g\in L^{2}(\Omega)$ and $T_{j}g=f$ Thus $0=\left\langle f\,,\,T_{j}g\right\rangle_{L^{2}(\mathbb{R})}=\|f\|_{L^{2}(\mathbb{R})}^{2}>0$, a contradiction . ∎
We want to show that $\Omega$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$. We will show next that $\Omega$ is contained in a larger set $\tilde{\Omega}$ which tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$.
The set $\Omega^{\prime}=\Omega+p\mathbb{Z}$ is clearly invariant under translations by $p\mathbb{Z}$ and we show that
(2.7) $\{\Omega^{\prime}+j:j\in\{0,1,\dots,p-1\}\}\mbox{ are disjoint sets}.$
If not, there exist $k_{1},k_{2}\in\mathbb{Z}$, $j_{1},j_{2}\in\{0,\dots,p-1\}$ such that $(\Omega+pk_{1}+j_{1})\cap(\Omega+pk_{2}+j_{2})\neq\emptyset$. We can assume $j_{1}>j_{2}$ and then $j_{1}-j_{2}\in\{0,\dots,p-1\}$ and $(\Omega+p(k_{1}-k_{2})+(j_{1}-j_{2}))\cap\Omega\neq\emptyset$, which contradicts Proposition 2.12.
Lemma 2.13.
Define
$\Omega_{0}:=\Omega^{\prime}\cap[0,p).$
  1. (i)
    The sets $\{\Omega_{0}+k:k\in\mathbb{Z}\}$ are disjoint.
  2. (ii)
    $\Omega^{\prime}=\Omega+p\mathbb{Z}=\Omega_{0}+p\mathbb{Z}$.
Proof.
Consider two sets $\Omega_{0}+pl+j$ and $\Omega_{0}+pl^{\prime}+j^{\prime}$ with $l,l^{\prime}\in\mathbb{Z}$ and $j,j^{\prime}\in\{0,\dots,p-1\}$. If $j\neq j^{\prime}$ then the intersection of these two sets is contained in $(\Omega^{\prime}+j)\cap(\Omega^{\prime}+j^{\prime})=\emptyset$. If $j=j^{\prime}$ and the intersection of the two sets is nonempty, then $(\Omega_{0}+pl)\cap(\Omega_{0}+pl^{\prime})\neq\emptyset$ and this is impossible because $\Omega_{0}$ is contained in $[0,p)$. This proves (i).
For (ii), we clearly have $\Omega_{0}+p\mathbb{Z}\subset\Omega+p\mathbb{Z}$. For the reverse inclusion, let $x\in\Omega^{\prime}$. Then, there exists $k\in\mathbb{Z}$ such that $x-pk\in[0,p)$. Since $\Omega^{\prime}$ is invariant under translations by $p\mathbb{Z}$, it follows that $x-pk\in\Omega^{\prime}\cap[0,p)=\Omega_{0}$ and thus $x\in\Omega_{0}+pk$. ∎
Lemma 2.14.
Define $R:=[0,1)\setminus((\Omega_{0}+\mathbb{Z})\cap[0,1))$. $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $\mathbb{Z}$ and $\tilde{\Omega}:=(\Omega_{0}\cup R)+p\mathbb{Z}$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$ and contains $\Omega$.
Proof.
Note that $\Omega_{0}+\mathbb{Z}$ and $R+\mathbb{Z}$ are disjoint. Otherwise, there exist $x_{0}\in\Omega_{0}$, $k\in\mathbb{Z}$, $r\in R$ and $l\in\mathbb{Z}$ such that $x_{0}+k=r+l$. So $r=x_{0}+(k-l)\in(\Omega_{0}+\mathbb{Z})\cap[0,1)$, which contradicts the definition of $R$. The sets $\{R+k:k\in\mathbb{Z}\}$ are disjoint because $R$ is contained in $[0,1)$. The sets $\{\Omega_{0}+k:k\in\mathbb{Z}\}$ are disjoint, by Lemma 2.13. It follows that the sets $\{(\Omega_{0}\cup R)+k:k\in\mathbb{Z}\}$ are disjoint. Also,
$(\Omega_{0}\cup R)+\mathbb{Z}=(\Omega_{0}+\mathbb{Z})\cup(R+\mathbb{Z})=((\Omega_{0}+\mathbb{Z})\cup R)+\mathbb{Z}\supseteq[0,1)+\mathbb{Z}=\mathbb{R}.$
This shows that $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $\mathbb{Z}$.
For $\tilde{\Omega}=(\Omega_{0}\cup R)+p\mathbb{Z}$, since $\Omega_{0}\cup R$ tiles $\mathbb{R}$ with $\mathbb{Z}$, we get that the sets $\{\tilde{\Omega}+j=(\Omega_{0}\cup R)+j+p\mathbb{Z}:j\in\{0,\dots,p-1\}\}$ are disjoint, and
$\tilde{\Omega}+\{0,\dots,p-1\}=(\Omega_{0}\cup R)+p\mathbb{Z}+\{0,\dots,p-1\}=(\Omega_{0}\cup R)+\mathbb{Z}=\mathbb{R}.$
Thus, $\tilde{\Omega}$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$.
By Lemma 2.13, $\Omega\subseteq\Omega_{0}+p\mathbb{Z}\subseteq\tilde{\Omega}$.
Lemma 2.15.
$\Omega=\tilde{\Omega}$ and therefore $\Omega$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$.
Proof.
Suppose the inclusion $\Omega\subset\tilde{\Omega}$ is proper, so the complement $\tilde{\Omega}\setminus\Omega$ has positive Lebesgue measure.
[…]
$L^{2}(\tilde{\Omega})\ni f\to\hat{f}|_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}=:\mathcal{F}_{r}f\in L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},p\,dx\right)$
is a surjective isometry. Therefore $\mathcal{F}_{r}f\perp\mathcal{F}_{r}g$, and also $\mathcal{F}_{r}g\neq 0$.
On the other hand, by hypothesis, $\mathcal{F}_{r}$ is a surjective isometry from $L^{2}(\Omega)$ to the same range $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},p\,dx\right)$. Thus $\{\mathcal{F}_{r}f:f\in L^{2}(\Omega)\}$ is the entire range $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},p\,dx\right)$; therefore, $\mathcal{F}_{r}g$ is orthogonal to the entire range, and thus it is orthogonal to itself, so it has to be zero, a contradiction.
In conclusion, $\Omega=\tilde{\Omega}$ and thus $\Omega$ tiles $\mathbb{R}$ with $\{0,\dots,p-1\}$, and this shows the converse implication in our Theorem 1.8. ∎
Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data Availability Statement. No new data were created or analyzed in this study. Data sharing is not applicable to this article.
Funding. No funds, grants, or other support were received during the preparation of this manuscript.
Author Contribution. All authors contributed equally to the study conception, design, data collection, analysis, and manuscript preparation. Both authors read and approved the final manuscript.